a woman is 5 times older than her youngest daughter. how old is each now if the sum of their ages is 50?
Answer by fcabanski(1391)

Call the daughter's age x and the woman (5 times older) 5x.


The sum of their ages is 50 so x + 5x = 50
6x = 50
x = 8 and 1/3


1/3 of a year is 4 months.


The daughter is 8 years 4 months old.


Mom is 5*(8 and 1/3) = 41 and 2/3 = 41 years 8 months.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com